# Phase error rate estimation with basis-efficiency mismatch for decoy-state BB84

### v1

@@ Abstract
We present a finite-size security proof of the BB84 QKD protocol against coherent attacks, using entropic uncertainty relations, in the presence of imperfectly characterized basis-efficiency mismatch. Our proof requires no new modifications to the protocol steps or hardware, and works by obtaining a suitable bound on the phase error rate. It is applicable to imperfectly characterized detectors, and only requires the maximum relative difference in detection efficiencies and dark count rates of the detectors to be characterized. Moreover, our proof allows Eve to use any spatio-temporal degree of freedom in each round to choose detector efficiencies and dark count rates in their allowed ranges, thereby addressing an important problem of detector side channels. We prove security in the variable-length framework, where users are allowed to adaptively determine the length of key to be produced, and number of bits to be used for error-correction, based on observations made during the protocol. We quantitatively demonstrate the effect of basis-efficiency mismatch by applying our results to the decoy-state BB84 protocol.
@@ I Introduction
Security proofs of QKD based on the entropic uncertainty relations (EUR) [1, 2, 3], and the phase error correction approach [4, 5], yield some of the highest key rates against coherent attacks in the finite-size regime. While source imperfections [6, 7, 8, 9] have been extensively studied within these the phase error correction framework, detector imperfections have not yet been addressed in any meaningful sense in either security proof framework. In particular, these proof techniques require the probability of detection in Bob’s detection setup to be independent of basis choice. This assumption is referred to as “basis-independent loss” in the literature, while the violation of this assumption is referred to as “basis-efficiency mismatch” or “detection-efficiency mismatch”. Satisfying this assumption requires the efficiency and dark count rates of Bob’s detectors to be exactly identical. Therefore, justifying this assumption in practice requires exact characterization of Bob’s identical detectors. Either of these tasks are impossible in practice. Therefore, these proof techniques are not applicable to practical prepare-and-measure (and entangled-based) QKD scenarios involving realistic detectors. This has led to many works advocating for the use of MDI-QKD [10] to address detector imperfections. To date, there is no security proof method that addresses this problem without requiring significant protocol modifications (such as MDI-QKD). While some theoretical analyses of basis-efficiency mismatch in the IID asymptotic setting exist for standard QKD [11, 12, 13, 14, 15], there is no work that handles basis-efficiency mismatch for coherent attacks in the finite-size regime. Meanwhile, there have been several experimental demonstrations [16, 17, 18, 19, 20], exploiting basis-efficiency mismatch for attacks on QKD systems.
In this work, we prove the security of the decoy-state BB84 protocol with an active detection setup without assuming basis-independent loss. We do so by showing that the phase error rate on the key generation can be suitably bounded even without the assumption. We explicitly define metrics $\delta_{1}$, $\delta_{2}$ that quantify the deviation from the ideal case, and bound the phase error rate in terms of these deviations. These metrics can be directly computed from the experimental characterization of the detectors. Our results extend the security of QKD to the following practical scenarios:
@@
Moreover, we prove security in the variable-length framework [23] which allows Alice and Bob to adaptively determine the number of bits to be used for error-correction, and the length of the output key, based on the observations during runtime. Such protocols are critical for practical implementations, where the honest behaviour of the channel connecting Alice and Bob is difficult to determine in advance. A security proof for variable-length QKD has been recently obtained for generic protocols in Ref. [24] which utilizes an IID security proof followed by the postselection technique lift [25, 26] to coherent attacks, in contrast to the EUR approach used in this work. The use of the postselection technique in Ref. [24] leads to pessimistic key rates, which is avoided by this work. Our proof of variable-length security builds upon prior work on variable-length security for qubit BB84 using the phase error correction approach by Tsurumaru and Hayashi [27]. We extend their approach by allowing for lossy protocols and decoy-state protocols, and also do not require additional assumptions made in Ref. [27] (such as perfect uniformity of Alice’s raw key).
This paper is organized as follows. In Section II, we describe the QKD protocol that we consider in this work. For simplicity, we first consider the BB84 protocol where Alice prepares perfect single-photon signal states. We do this since there are many existing techniques for dealing with imperfect state preparation by Alice, and the goal of this work is to focus on detector imperfections. In Section III we prove that the variable-length security of the QKD protocol follows if one is able to obtain suitable bounds on the phase error rate (Eq. 7). In Section IV we show how such bounds can be obtained in the case where the basis-independent loss assumption is satisfied. In Section V, we show how such bounds can be obtained in scenarios where the basis-independent loss assumption is not satisfied. We also discuss similarities and differences with other related work [11, 12, 15, 14, 13] towards the end of this section. In Section VI we extend our analysis to prove the variable-length security of the decoy-state BB84 protocol with imperfect detectors. In Section VII we apply our results and compute key rates for the decoy-state BB84 protocol, and study the impact of basis-efficiency mismatch on key rates. We base our analysis on Lim et al [28], and also point out and fix a few technical issues in that work. In Section VIII we explain how our results can be applied to detector side channels (Case. 3 above), and generalize to independent detectors. In Section IX we summarize and present concluding remarks. Many details are relegated to the Appendices.
@@
    Provide a method for phase error rate estimation in the presence of imperfectly characterized basis-efficiency mismatch, in the finite-site setting against coherent attacks.
@@
  3. 3.
    Resolve detector side-channel vulnerabilities by allowing Eve to induce mismatch in detectors through the use of spatio-temporal modes.
  4. 4.
    Address the case where Bob’s detectors are independent (but not IID).
We have attempted to write this paper in a largely modular fashion beyond Section II. For variable-length security, one can directly read Sections III and VI.5. Similarly, for phase error estimation for the BB84 protocol, one can directly read Sections IV and V. For phase error estimation in decoy-state BB84, one can read Sections VI.2 and VI.3. For a recipe for applying our results to compute key rates in the presence of basis efficiency mismatch, one can read Section VII.2.
@@ II Protocol Description
    Now, if Alice prepares the states from Eq. 1, her POVM elements corresponding to the basis $b$ signal states sum to $p^{(A)}_{(b)}\mathrm{I}_{A}$. Because of this fact, can one view Alice’s measurement process, after using the source-replacement scheme, as equivalent to choosing the basis $Z(X)$ with probability $p^{(A)}_{(Z)}(p^{(A)}_{(X)})$, followed by measuring using the POVM $\{\Gamma^{(A)}_{(b,0)},\Gamma^{(A)}_{(b,1)}\}$, for a given basis $b$. This reflects the fact that Eve has no knowledge of the basis used.
@@
    Without loss of generality, one can always use the source-replacement scheme, and delay Alice’s measurements until after Eve’s attack has been completed, for any set of signal states. However, this process might result in POVM elements for Alice whose sum (for a specific basis) is not proportional to identity. In this case, Alice’s measurements are incompatible with active basis choice after the source-replacement scheme. Active basis choice for Alice is required when using the EUR statement (Section III), and in bounding the phase error rate (Sections IV and V). It is precisely for this reason that Eq. 1 is needed. For methods to address imperfect state preparation, we refer the reader to [8, 7, 6, 9] (however we note that the analysis there is within the phase error framework).
@@
    where Alice’s POVMs are defined in Eq. 2, and Bob’s in Section VII.1.
@@
    As we will see in Section IV, the mathematical assumption on Bob’s detector setup needed for phase error estimation is actually given by
@@
    This means that the probability of a round being inconclusive (i.e discarded) is independent of the basis. Notice that Eq. 4 depends on the choice of classical post-processing on Bob’s side. In particular, it can be trivially satisfied by mapping no-click and double-click events to 00 and $1$ randomly (so that $\Gamma^{(B)}_{(X,\bot)}=\Gamma^{(B)}_{(Z,\bot)}$ is zero). However, such a protocol cannot produce a key when loss is greater than $50\%$, and is therefore impractical. In general, if one assumes the canonical model of detectors (see Section VII.1), and maps double-clicks to 0/1 randomly, then Eq. 4 requires the loss and dark count rates in each detector-arm to be equal. This is why this condition is referred to as “basis-independent loss”, and its violation is referred to as “detection-efficiency mismatch” in the literature. Note that even for identical detectors, one is forced remap double-click events to satisfy Eq. 4.
@@
III Variable-length security proof
In this section we will prove the variable-length security of the protocol from Section II. In order to prove the $\varepsilon_{\text{secure}}$-security of a variable-length QKD protocol [23], one must show that for all attacks by the adversary, the following statement is true :
Requirements on phase error estimation procedure: Note that in a QKD protocol, one starts with a fixed but unknown state $\rho_{A^{n}B^{n}E^{n}}$ that represents Eve’s attack. This state then gives rise to random variables $\bm{n_{X}},\bm{n_{K}},\bm{e^{\text{obs}}_{X}},\bm{e^{\text{key}}_{X}}$. Here $\bm{e^{\text{key}}_{X}}$ denotes the random variable corresponding to the “phase error rate” in the key-generation rounds, when Alice and Bob measure those rounds (virtually) in the $X$ basis. (The phase error rate is explained in greater detail in Sections IV and V). To obtain security, one must obtain a high probability upper bound on the phase error rate $\bm{e^{\text{key}}_{X}}$. We assume that one has a way to obtain the following statement (which we prove in Sections IV and V):
@@ III Variable-length security proof
This states that the phase error rate is upper bounded (with high probability) by a suitable function $\mathcal{B}_{\delta_{1},\delta_{2}}$ of the observed error rate in the $X$ basis rounds, and the number of test and key generation rounds. The function $\mathcal{B}_{\delta_{1},\delta_{2}}$ depends on the metrics $\delta_{1},\delta_{2}$ that quantify the deviation from ideal behaviour for a given protocol description, and are computed in Section VII. Let
$\leq\varepsilon_{\text{AT}}^{2},$
where the sum is over all possible values of $n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z}$. (In this work, we will use the convention that $\sum_{x}$ denotes the sum over all possible values $x$ can take). We will utilize Eq. 9, which follows from Eq. 7, in proving the variable-length security of the QKD protocol in Theorem 1 below. We will obtain a suitable $\mathcal{B}_{\delta_{1},\delta_{2}}$ satisfying Eq. 7 in Sections IV and V, with and without the basis-independent loss assumption.
@@
Theorem 1 (Variable-length security of BB84 with qubit source).
Suppose Eq. 7 is satisfied and let $\lambda_{\text{EC}}(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ be a function that determines the number of bits used for error-correction. Define
@@
where $h(x)$ is the binary entropy function for $x\leq 1/2$, and $h(x)=1$ otherwise. Then the variable-length QKD protocol that produces a key of length $l(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ using $\lambda_{\text{EC}}(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ bits for error-correction, upon the event $\Omega_{(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})}\wedge\Omega_{\text{EV}}$ is $(2\varepsilon_{\text{AT}}+\varepsilon_{\text{PA}}+\varepsilon_{\text{EV}})$-secure 444 For pedagogical reasons, we ignore the issues arising from non-integer values of hash-lengths. Such issues can be easily fixed by suitable use of floor and ceiling functions..
Proof.
Our proof consists of two parts. In the first part we will use entropic incertainty relations and Eq. 7 to obtain a suitable lower bound on the smooth min entropy of the raw key register in the QKD protocol. In the second part, we use this to prove variable-length security.
First, focus on the state $\rho_{A^{n_{K}}B^{n_{K}}E^{n}C^{n}|\Omega_{(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})}}$, which is the state on the detected key generation rounds. This state can be obtained by transforming Bob’s measurement procedure to consist of two steps, and then only implementing the first step measurement which determines the detect vs no-detect outcome. Such a state can be rigorously obtained using Lemma 1 from Section IV. For the purposes of this proof, we only need the fact that it is well defined. We will obtain a bound on the smooth min entropy of the key generated from this state.
Remark 6.
We can make Bob measure his systems $B$ in the $X$ basis to obtain the classical outcome $X_{B}$. Then, using data processing [35, Theorem 6.2], we obtain
which along with Eq. 12 gives us
$\leq\varepsilon_{\text{PA}}+2\varepsilon_{\text{AT}}.$
Here, we used the leftover-hashing lemma [2, Proposition 9] with the appropriate smoothing parameter on the sub-normalized state $(\rho_{|\Omega_{(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})}})_{\wedge\Omega_{\text{EV}}}$ for the first inequality. Since we use sub-normalized conditioning on $\Omega_{\text{EV}}$, it only appears in the smooth min entropy term and not inside the probability. Next, we use [2, Lemma 10] to get rid of the sub-normalized conditioning ($\wedge\Omega_{\text{EV}}$) in the smooth min entropy term in the second inequality. We used [35, Lemma 6.8] to split off the error-correction information ($\lambda_{\text{EC}}(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$) and error-verification information ($\log(2/\varepsilon_{\text{EV}})$) in the third inequality. We used the bound on the smooth min entropy from Eq. 14 for the fourth inequality, and the values of $l(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ and $\lambda_{\text{EC}}(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ from Eq. 10 for the fifth equality. We used concavity of the square root function and Jensen’s inequality to pull the sum over probabilities inside the square root for the sixth inequality, and Eq. 9 for the final inequality.
Remark 7.
Note that the critical step here was using the concavity of the square root function to see that a bound on the average failure probability of the phase error estimation procedure is enough to prove security. This is the same fundamental trick used by Hayashi and Tsurumaru in their variable-length proof using the phase error correction approach, in Ref. [27]. Our presentation here is in the EUR framework, where this manifests in the deliberate choice of our smoothing parameter in the first part of the proof. We are also able to deal with loss, unlike Ref. [27], and does not require some additional assumptions (such as uniformity of Alice’s raw key) made in Ref. [27].
Remark 8.
It is interesting to note that these issues are completely avoided by the above proof, in a very different manner than [24]. This is because in this proof, we always apply the leftover-hashing lemma on a state conditioned on the specific length of the raw key register (Eq. 14). Therefore, the leftover-hashing lemma can be applied in a straightforward manner, and there are no issues is choosing the hashing family based on the specific length of the raw key register. In other words, the PA procedure described above is valid for this proof.
In a similar sense, the variable-length security proof from [24] critically relied on a technical lemma (Lemma 9), that necessitated the use of Rényi entropies instead of smooth min entropy in that work. However, the variable-length proof presented above takes a different approach, and does not impose the same requirements on the behaviour of smooth min entropy. Again, this is due to the use of Eq. 14.
Thus, we have proven the variable-length security of a QKD protocol, in the EUR framework, assuming that our phase error estimation procedure satisfies Eq. 7. We now obtain such a bound on the phase error in the next section.
@@
IV Phase error estimation with basis-independent loss assumption
[…]
Figure 1: Protocol flowchart for the equivalent protocol from Section IV, where basis-independent loss assumption (Eq. 4) is satisfied. The dotted arrows and boxes represent virtual measurements that do not actually happen in the real protocol. Connections between different boxes are highlighted using curved arrows. We use the Serfling bound (Lemma 2) to obtain a bound on the phase error rate from observations. The phase error rate is then used to bound the smooth min entropy using the EUR statement. We use $\bm{N}_{P}$ to denote the number of $P$ measurement outcomes, where $P$ denotes a POVM element. For clarity, we have omitted the conditioning on events in the figure (but not in our proof). The basis used for measurements is indicated in each box, and refers to the basis used by both Alice and Bob.

### v2

# Phase error rate estimation in QKD with imperfect detectors

@@ Abstract
We present a finite-size security proof of the decoy-state BB84 QKD protocol against coherent attacks, using entropic uncertainty relations, for imperfect detectors. We apply this result to the case of detectors with imperfectly characterized basis-efficiency mismatch. Our proof works by obtaining a suitable bound on the phase error rate, without requiring any new modifications to the protocol steps or hardware. It is applicable to imperfectly characterized detectors, and only requires the maximum relative difference in detection efficiencies and dark count rates of the detectors to be characterized. Moreover, our proof allows Eve to choose detector efficiencies and dark count rates in their allowed ranges in each round, thereby addressing an important problem of detector side channels. We prove security in the variable-length framework, where users are allowed to adaptively determine the length of key to be produced, and number of bits to be used for error-correction, based on observations made during the protocol. We quantitatively demonstrate the effect of basis-efficiency mismatch by applying our results to the decoy-state BB84 protocol.
@@ I Introduction
Security proofs of QKD based on the entropic uncertainty relations (EUR) [1, 2, 3], and the phase error correction approach [4, 5], yield some of the highest key rates against coherent attacks in the finite-size regime. While source imperfections [6, 7, 8, 9] have been extensively studied within the phase error correction framework, detector imperfections have not yet been addressed in any meaningful sense in either security proof framework. In particular, these proof techniques require the probability of detection in Bob’s detection setup to be independent of basis choice. This assumption is referred to as “basis-independent loss” in the literature, while the violation of this assumption is referred to as “basis-efficiency mismatch” or “detection-efficiency mismatch”. Satisfying this assumption requires the efficiency and dark count rates of Bob’s detectors to be exactly identical. Therefore, justifying this assumption in practice requires exact characterization of Bob’s identical detectors. Either of these tasks are impossible in practice. Therefore, these proof techniques are not applicable to practical prepare-and-measure (and entangled-based) QKD scenarios involving realistic detectors. This has led to many works advocating for the use of MDI-QKD [10] to address detector imperfections. To date, there is no security proof method that addresses this problem without requiring significant protocol modifications (such as MDI-QKD). While some theoretical analyses of basis-efficiency mismatch in the IID asymptotic setting exist for standard QKD [11, 12, 13, 14, 15], there is no work that handles basis-efficiency mismatch for coherent attacks in the finite-size regime. Meanwhile, there have been several experimental demonstrations [16, 17, 18, 19, 20], exploiting basis-efficiency mismatch for attacks on QKD systems.
In this work, we prove the security of the decoy-state BB84 protocol with an active detection setup without assuming basis-independent loss. We do so by showing that the phase error rate can be suitably bounded even without the assumption. We explicitly define metrics $\delta_{1}$, $\delta_{2}$ that quantify the deviation from the ideal case, and bound the phase error rate in terms of these deviations. We compute these metrics for the case of detectors with basis-efficiency mismatch, directly from the experimental characterization of the detectors. Our results extend the security of QKD to the following practical scenarios:
@@
Moreover, we prove security in the variable-length framework [23] which allows Alice and Bob to adaptively determine the number of bits to be used for error-correction, and the length of the output key, based on the observations during runtime. Such protocols are critical for practical implementations, where the honest behaviour of the channel connecting Alice and Bob is difficult to determine in advance. A security proof for variable-length QKD has been recently obtained for generic protocols in Ref. [24] which utilizes an IID security proof followed by the postselection technique lift [25, 26] to coherent attacks. The use of the postselection technique in Ref. [24] leads to pessimistic key rates, which is avoided by this work. Our proof of variable-length security uses the same essential tricks as prior work on variable-length security in Refs. [27] and [28, Chapter 3] for the phase error correction framework, and is nearly identical to [29, Supplementary Note A] for the EUR framework.
This paper is organized as follows. In Section II, we describe the QKD protocol that we consider in this work. For simplicity, we first consider the BB84 protocol where Alice prepares perfect single-photon signal states. We do this since there are many existing techniques for dealing with imperfect state preparation by Alice, and the goal of this work is to focus on detector imperfections. We show that that the variable-length security of the QKD protocol follows if one is able to obtain suitable bounds on the phase error rate (Eq. 5). In Section III we show how such bounds can be obtained in the case where the basis-independent loss assumption is satisfied. In Section IV, we show how such bounds can be obtained in scenarios where the basis-independent loss assumption is not satisfied. We also discuss similarities and differences with other related work [11, 12, 15, 14, 13] towards the end of this section. In Section V we extend our analysis to prove the variable-length security of the decoy-state BB84 protocol with imperfect detectors. In Section VI we apply our results and compute key rates for the decoy-state BB84 protocol, and study the impact of basis-efficiency mismatch on key rates. We base our analysis of decoy-state BB84 on Lim et al [30], and also point out and fix a few technical issues in that work. In Section VII we explain how our results can be applied to detector side channels (Case. 3 above), and generalize to independent detectors. In Section VIII we summarize and present concluding remarks. Many details are relegated to the Appendices.
@@
    Provide a method for phase error rate estimation in the presence of (bounded) detector imperfections, in the finite-size setting against coherent attacks.
@@
    Resolve detector side-channel vulnerabilities by allowing Eve to control (bounded) detector imperfections.
  3. 3.
@@
We have attempted to write this paper in a largely modular fashion beyond Section II. For variable-length security, one can directly read Appendices B and V.5. Similarly, for phase error estimation for the BB84 protocol, one can directly read Section III (for perfect detectors) and Section IV (for imperfect detectors). For phase error estimation in decoy-state BB84, one can read Sections V.2 and V.3. For a recipe for applying our results to compute key rates in the presence of detector imperfections, one can refer to Section VI.1. For the application of our results to detectors with efficiency mismatch, one can refer to Sections VI.2, VI.3 and VI.4.
@@ II Protocol Description
    Now, if Alice prepares the states from Eq. 1, her POVM elements corresponding to the basis $b$ signal states sum to $p^{(A)}_{(b)}\mathrm{I}_{A}$. Because of this fact, one can view Alice’s measurement process, after using the source-replacement scheme, as equivalent to choosing the basis $Z(X)$ with probability $p^{(A)}_{(Z)}(p^{(A)}_{(X)})$, followed by measuring using the POVM $\{\Gamma^{(A)}_{(b,0)},\Gamma^{(A)}_{(b,1)}\}$, for a given basis $b$. This reflects the fact that Eve has no knowledge of the basis used.
@@
    Without loss of generality, one can always use the source-replacement scheme, and delay Alice’s measurements until after Eve’s attack has been completed, for any set of signal states. However, this process might result in POVM elements for Alice whose sum (for a specific basis) is not proportional to identity. In this case, Alice’s measurements are incompatible with active basis choice after the source-replacement scheme. We utilize the fact that Alice implements active basis choice when using the EUR statement (Appendix B), and in bounding the phase error rate (Sections III and IV). It is precisely for this reason that Eq. 1 is needed. For methods to address imperfect state preparation, we refer the reader to [8, 7, 6, 9] (however we note that the analysis there is within the phase error framework).
@@
    where Alice’s POVMs are defined in Eq. 2, and Bob’s in Section VI.2.
@@
    As we will see in Section III, the mathematical assumption on Bob’s detector setup needed for phase error estimation is actually given by
@@
    This means that the probability of a round being inconclusive (i.e discarded) is independent of the basis. Notice that Eq. 4 depends on the choice of classical post-processing on Bob’s side. In particular, it can be trivially satisfied by mapping no-click and double-click events to 00 and $1$ randomly (so that $\Gamma^{(B)}_{(X,\bot)}=\Gamma^{(B)}_{(Z,\bot)}$ is zero). However, such a protocol cannot produce a key when loss is greater than $50\%$, and is therefore impractical. In general, if one assumes the canonical model of detectors (see Section VI.2), and maps double-clicks to 0/1 randomly, then Eq. 4 requires the loss and dark count rates in each detector-arm to be equal. This is why this condition is referred to as “basis-independent loss”, and its violation is referred to as “detection-efficiency mismatch” in the literature. Note that even for identical detectors, one is forced remap double-click events to satisfy Eq. 4.
@@
II.1 Requirements on phase error estimation procedure and variable-length security
We now turn our attention to the estimation of the phase error rate. Note that in a QKD protocol, one starts with a fixed but unknown state $\rho_{A^{n}B^{n}E^{n}}$ that represents Eve’s attack. This state then gives rise to random variables $\bm{n_{X}},\bm{n_{K}},\bm{e^{\text{obs}}_{X}},\bm{e^{\text{key}}_{X}}$. Here $\bm{e^{\text{key}}_{X}}$ denotes the random variable corresponding to the “phase error rate” in the key-generation rounds, when Alice and Bob measure those rounds (virtually) in the $X$ basis. (The phase error rate is explained in greater detail in Sections III and IV). To obtain variable-length security, one must obtain a high probability upper bound on the phase error rate $\bm{e^{\text{key}}_{X}}$. We assume that one has a way to obtain the following statement (which we prove in Sections III and IV):
@@ II.1 Requirements on phase error estimation procedure and variable-length security
This states that the phase error rate is upper bounded (with high probability) by a suitable function $\mathcal{B}_{\delta_{1},\delta_{2}}$ of the observed error rate in the $X$ basis rounds, and the number of test and key generation rounds. We will obtain a suitable $\mathcal{B}_{\delta_{1},\delta_{2}}$ satisfying Eq. 5 in Sections III and IV, with and without the basis-independent loss assumption. The function $\mathcal{B}_{\delta_{1},\delta_{2}}$ depends on the metrics $\delta_{1},\delta_{2}$ that quantify the deviation from ideal behaviour for a given protocol description. We compute explicit bounds for $\delta_{1},\delta_{2}$ for detectors with efficiency mismatch in Section VI.3 using the recipe outlined in Section VI.1.
@@
Given an upper bound on the phase error rate (Eq. 5), we have the following theorem regarding the variable-length security of the QKD protocol described above. The proof is essentially identical to [29, Supplementary Note A], and uses the same techniques as those in Refs. [27, 28] and is included in Appendix B.
Theorem 1.
[Variable-length security of BB84 with qubit source] Suppose Eq. 5 is satisfied and let $\lambda_{\text{EC}}(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ be a function that determines the number of bits used for error-correction. Define
@@
where $h(x)$ is the binary entropy function for $x\leq 1/2$, and $h(x)=1$ otherwise. Then the variable-length QKD protocol that produces a key of length $l(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ using $\lambda_{\text{EC}}(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})$ bits for error-correction, upon the event $\Omega_{(n_{X},n_{K},e^{\text{obs}}_{X},e^{\text{obs}}_{Z})}\wedge\Omega_{\text{EV}}$ is $(2\varepsilon_{\text{AT}}+\varepsilon_{\text{PA}}+\varepsilon_{\text{EV}})$-secure 333 For pedagogical reasons, we ignore the issues arising from non-integer values of hash-lengths. Such issues can be easily fixed by suitable use of floor and ceiling functions..
@@
III Phase error estimation with basis-independent loss assumption